OpenStudy (hellokitty17):

What are the approximate solutions of 2x2 + x = 14, rounded to the nearest hundredth? x ≈ -2.41 and x ≈ 2.91 x ≈ -2.91 and x ≈ 2.41 x ≈ -5.82 and x ≈ 4.82 No real solutions

OpenStudy (hellokitty17):

@Mehek14

oregonduck (oregonduck):

Bring 14 to the left: 2x² + x - 14 = 0 The quadratic formula is x = (-b ± √(b² - 4ac))/2a. The quadratic equation is ax² + bx + c = 0. Furthermore, by using the formulas here, we have: a = 2 b = 1 c = -14 x = (-1 ± √(1² - 4(2)(-14)))/2(2) x = (-1 ± √(1 + 112))/4 x = (-1 ± √(113))/4 By calculator, we have: x ≈ 2.41 and x ≈ -2.91

oregonduck (oregonduck):

so b

oregonduck (oregonduck):

right @Mehek14 ????????

Mehek (mehek14):

yea i was just checking to see if it's correct

oregonduck (oregonduck):

ok